OpenStudy (anonymous):

The slope of the line tangent to the curve y^2+(xy+1)^3=0 at (2,-1) is what?

OpenStudy (anonymous):

The slope of the tangent line is given by \(\dfrac{dy}{dx}\). Finding that requires implicit differentiation: \[y^2+(xy+1)^3=0\] Differentiate both sides with respect to \(x\). \[2y\frac{dy}{dx}+3(xy+1)^2\left(y+x\frac{dy}{dx}\right)=0\] Now just solve for \(\dfrac{dy}{dx}\), then plug in \(x=2\) and \(y=-1\).